OpenStudy (anonymous):

Choose the equation of the vertical line that passes through the point (2, 4). A)x=4 B)x=2 C)y=2 D)y=4

OpenStudy (ivettef365):

vertical line means the slope = 0

OpenStudy (ivettef365):

therefore y = mx+b where m=0, substitute the x = 2 and y = 4 then you have 4=(0)(2)+b b=4

OpenStudy (ivettef365):

that means the y-int=4